OpenStudy (janaelovexoxo):

Which value is equivalent to 7.2 kilograms? 720 grams 7,200 grams 72,000 grams 720 milligrams 7,200 milligrams

OpenStudy (hoot_hoot):

1 000 g = 1 kg 1 000 000 mg = 1 kg \[ 7.2 kg = \frac{ 1 000 g }{ 1 kg } = \frac{ 1 000 000 mg }{ 1 kg }\] 1 kg is at the denominator to be able to cancel out the unit kg so you'll be left with g (grams) or mg (milligrams). Following the formula, you need to multiply 7.2 by either 1 000 (grams) or 1 000 000 (milligrams). What'd you get?

OpenStudy (hoot_hoot):

Sorry. It should be times not equal. \[7.2kg \times \frac{ 1 000 g}{ 1 kg } = ?\] OR \[7.2kg \times \frac{ 1 000 000 mg}{ 1 kg } = ?\]